Find the measure of an angle whose supplement is 10 degrees more than twice its complement. Let 90 - x equal the degree measure of its complement and 180-x equal the degree of measure of its supplement. how do I write the equation?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(48528)

Find the measure of an angle whose supplement is 10 degrees more than twice its complement. Let 90 - x equal the degree measure of its complement and 180-x equal the degree of measure of its supplement. how do I write the equation?
----------------------------------------------
Let the angle be "x"
Supplement is "180-x"
Complement is "90-x"
--------------------------
Equation:
supplement = 2*complement + 10
180-x = 2(90-x) + 10
180-x = 180 - 2x + 10
-x = -2x+10
x = 10 degrees
==================================
Cheers,
Stan H.

Answer by solver91311(12126)



Just like it is written in words:

An angle whose supplement () is (=) 10 more than (10 + ) twice its complement ().




John